OpenStudy (anonymous):

Suppose f(x) = A ln(x) + B, where A and B are constants. If f(1) = 8 and f(e) = 5, what are A and B?

OpenStudy (anonymous):

i can post the answers too if you want

OpenStudy (turingtest):

f(1)=A ln(1)+B=A(0)+B=B=8 f(e)=A ln(e)+8=A+8=5--->A=-3

OpenStudy (anonymous):

A ln (1) + B = 8 A ln (e) + B = 5 by definition ln e = 1 and ln 1 = 0 so A + B = 5 and B = 8 so A = -3

OpenStudy (anonymous):

What do I press on my calc for In?